# Infinite-dimensional next-generation reservoir computing

### v2

@@ I Introduction
Despite the ease of training associated with the RC methodology, its implementation depends heavily on hyperparameters that are sometimes difficult to estimate robustly. This difficulty has motivated various authors to replace the standard RC approach with nonlinear vector autoregression in which the covariates are monomials constructed using the previous inputs. In these methods, the only hyperparameters to be chosen are the maximum order of the monomials and the number of lags of past signals. This approach has been called next-generation reservoir computing (NG-RC) [13, 5, 4, 26, 25, 42] and has displayed excellent performances in control and spatio-temporal forecasting tasks.
@@
The last part of the paper contains various numerical simulations that illustrate that (i) using the more computationally tractable polynomial kernel regression, one has access to a broader feature space, which allows learning more complex systems than those typically presented when using the NG-RC methodology and (ii) the Volterra kernel is a useful tool that can produce more accurate forecasts because it allows one to take infinite lags and monomial orders into account, which is of relevance in the modeling of long-memory phenomena that exhibit functionally complex dependencies. All the codes and data necessary to reproduce the results in the paper are available at https://github.com/Learning-of-Dynamic-Processes/kernelngrcvolterra.
@@ II.1 Notation
We often work with input or output sequence spaces. Let $\mathbb{Z}$ denote the set of integers and ${\mathbb{Z}_{-}}$ be the set of non-positive integers. Denote by $\mathbb{R}$ the set of real numbers and the $d$-dimensional Euclidean space by $\mathbb{R}^{d}$. Let the space of sequences of $d$-dimensional real vectors indexed by ${\mathbb{Z}_{-}}$ be denoted $\left(\mathbb{R}^{d}\right)^{\mathbb{Z}_{-}}$. Denote generically the input space by $\mathcal{Z}$ and the output space by $\mathcal{Y}$, which, in most cases, will be subsets of a finite-dimensional Euclidean space. Given $\tau\in\mathbb{N}$, we shall denote by $\left(\mathbb{R}^{d}\right)^{\tau}$ the space of $\tau$-long sequences of $d$-dimensional real inputs indexed by $\{-\tau+1,\ldots,-1,0\}$. Given the input space $\mathcal{Z}\subset\mathbb{R}^{d}$, we will denote the set of semi-infinite sequences with elements in ${\cal Z}$ by ${\cal Z}^{\mathbb{Z}_{-}}$. The components of the sequences $\underline{\mathbf{z}}\in{\cal Z}^{\mathbb{Z}_{-}}$ are given by ${\bf z}_{i}\in\mathbb{R}^{d}$, $i\in\mathbb{Z}_{-}$, that is, $\underline{\mathbf{z}}=\left({\bf z}_{i}\right)_{i\in\mathbb{Z}_{-}}$. Now, given $i\in\mathbb{Z}_{-}$, $\tau\in\mathbb{N}$, and $\underline{\mathbf{z}}\in{\cal Z}^{\mathbb{Z}_{-}}$, we denote by $\underline{\mathbf{z}_{i}}\coloneqq(\ldots,\mathbf{z}_{i-1},\mathbf{z}_{i})\in{\cal Z}^{\mathbb{Z}_{-}}$ and by $\underline{\mathbf{z}_{i}}^{\tau}\coloneqq(\mathbf{z}_{i-\tau+1},\ldots,\mathbf{z}_{i})\in{\cal Z}^{\tau}$.
@@ II.2 Data generating processes
Throughout this paper, we are interested in the learning and forecasting of discrete-time, time-invariant, and causal dynamic processes that are generated by functionals of the form $H:\left(\mathbb{R}^{d}\right)^{\mathbb{Z}_{-}}\longrightarrow\mathcal{Y}$, that is, given $\underline{{\bf z}}\in({\mathbb{R}}^{d})^{\mathbb{Z}}$, we have that ${\bf y}_{t}\in H(\underline{{\bf z}_{t}})$, $t\in\mathbb{Z}$. In practice, the inputs $\underline{{\bf z}}\in({\mathbb{R}}^{d})^{\mathbb{Z}}$ can be just deterministic sequences or realizations of a stochastic process. A related data generating process that we shall be using are the causal chains with infinite memory (CCIM) [10, 11, 1]. These are infinite sequences $(\mathbf{y}_{i})_{i\in\mathbb{Z}}$, where $\mathbf{y}_{i}\in\mathcal{Y}$ for all $i\in\mathbb{Z}$, are such that $\mathbf{y}_{i}=H(\underline{\mathbf{y}_{i-1}})$ for all $i\in\mathbb{Z}$ and some functional $H:{\cal Y}^{\mathbb{Z}_{-}}\rightarrow{\cal Y}$. Takens’ Theorem [45] and its generalizations [27, 28, 18, 19] guarantee that the low-dimensional observations of dynamical systems follow, under certain conditions, a dynamical prescription of this type.
@@ II.3 Kernel methods
Kernels and feature maps. A kernel on $\mathcal{Z}$ is a function $K:\mathcal{Z}\times\mathcal{Z}\longrightarrow\mathbb{R}$ that is symmetric and positive semi-definite. In this context, positive semi-definiteness means that for any $\alpha_{i},\alpha_{j}\in\mathbb{R}$, $\mathbf{z}_{i},\mathbf{z}_{j}\in\mathcal{Z}$, $i,j\in\{1,\ldots,n\}$,
@@
Let $\mathbb{H}$ be a Hilbert space of real-valued functions on $\mathcal{Z}$ endowed with pointwise sum and scalar multiplication. Denote the inner product on $\mathbb{H}$ by $\langle\cdot,\cdot\rangle_{\mathbb{H}}$. We say that $\mathbb{H}$ is a reproducing kernel Hilbert space (RKHS) associated to the kernel $K$ if the following two conditions hold. First, for all $\mathbf{z}\in\mathcal{Z}$, we have that the functions $K(\cdot,\mathbf{z})\in\mathbb{H}$, and for all $\mathbf{z}\in\mathcal{Z}$ and all $f\in\mathbb{H}$, the reproducing property, $f(\mathbf{z})=\langle f,K(\cdot,\mathbf{z})\rangle_{\mathbb{H}}$, $z\in{\cal Z}$, is satisfied. The maps of the form $K_{\mathbf{z}}(\cdot)\coloneqq K(\cdot,\mathbf{z}):\mathcal{Z}\longrightarrow\mathbb{R}$ are called kernel sections. Second, the Dirac functionals $\delta_{{\bf z}}:\mathbb{H}\rightarrow\mathbb{R}$ defined by $\delta_{{\bf z}}(f):=f({\bf z})$ are continuous, for all ${\bf z}\in{\cal Z}$.
In this setup one may define the map $\Phi_{\mathbb{H}}:\mathcal{Z}\longrightarrow\mathbb{H}$ given by $\Phi_{\mathbb{H}}(\mathbf{z})\coloneqq K_{\mathbf{z}}$, ${\bf z}\in{\cal Z}$. Then from the reproducing property of the RKHS, the kernel function can be written as the inner product of kernel sections. Indeed, given ${\bf z},{\bf z}^{\prime}\in{\cal Z}$. we can write
@@
where the bar denotes the completion with respect to the metric induced by the inner product $\langle\cdot,\cdot\rangle_{\mathbb{H}}$. This inner product obviously satisfies that for $f=\sum^{n}_{i=1}\alpha_{i}K_{\mathbf{z}_{i}},g=\sum^{m}_{j=1}\beta_{j}K_{\mathbf{z}_{j}}\in\mathbb{H}$, $\langle f,g\rangle_{\mathbb{H}}\coloneqq\sum_{i=1}^{n}\sum_{j=1}^{m}\alpha_{i},\beta_{j}K(\mathbf{z}_{i},\mathbf{z_{j}})$. We use the symbol $\|\cdot\|_{\mathbb{H}}$ for the norm induced by $\langle\cdot,\cdot\rangle_{\mathbb{H}}$.
The construction we just presented produces an RKHS and a feature map out of a kernel map. Conversely, given any Hilbert space $(H,\left\langle\cdot,\cdot\right\rangle_{H})$ and a map $\Phi:\mathcal{Z}\rightarrow H$ from the set ${\cal Z}$ into $H$, we can construct a kernel map that has $\Phi$ as a feature map and $H$ as a feature space, that is, define
@@
Such a function $K$ is clearly symmetric and positive semi-definite; thus, it satisfies the conditions needed to be a kernel function. It can actually be proved (see [7, Theorem 4.16]) that a map $K:\mathcal{Z}\times\mathcal{Z}\longrightarrow\mathbb{R}$ is a kernel if and only if there exists a feature map that allows $K$ to be represented as in (2). Note that given a kernel function, the feature representation (2) is not unique in the sense that neither the feature space $H$ nor the feature map $\Phi:\mathcal{Z}\rightarrow H$ are unique. In particular, given $K$, the canonical feature map and the RKHS are a choice of feature map and feature space for $K$, respectively.
@@
=$\mathbb{H}=\{f:\mathcal{Z}\rightarrow\mathbb{R}\mid f(\mathbf{z})=\left<w,\Phi(\mathbf{z})\right>_{H},w\in H,\mathbf{z}\in\mathcal{Z}\},$ (3)
@@
[…]
This time around, the bar denotes the uniform closure. Note that the continuity of $K$, the reproducing property, and the compactness of ${\mathcal{K}}$ imply that the uniform closure that defines $K\left({\mathcal{K}}\right)$ contains the completion of the vector space ${\rm span}\left\{K_{z}\mid z\in{\cal Z}\right\}$. The continuous kernel $K$ is called kernel universal when for any compact subset $\mathcal{K}\subset\mathcal{Z}$,
$K(\mathcal{K})=C^{0}(\mathcal{K}).$
Many kernels used in practice are universal, e.g., the Gaussian kernel on Euclidean space.
Kernel and nonlinear regressions. The universal approximation properties of RKHSs that we just described make them good candidates to serve as concept spaces in nonlinear regressions. More explicitly, suppose that ${\cal Z}\subset{\mathbb{R}}^{d}$, $\mathcal{Y}\subset\mathbb{R}$, and that a function $f:{\cal Z}\rightarrow{\cal Y}$ needs to be estimated using a finite sample of $n$ pairs of input and outputs $\{(\mathbf{z}_{i},y_{i}:=f(\mathbf{z}_{i}))\}_{i=1}^{n}$ where $\mathbf{z}_{i}\in\mathcal{Z}$ and $y_{i}\in\mathcal{Y}$, $i=1,\ldots,n$. If the estimation is carried out using a empirical risk with squared loss
@@
and the hypothesis set is the RKHS $\mathbb{H}$ corresponding to a kernel $K$, the Representer Theorem states that the minimizer of the ridge regularized empirical risk on $\mathbb{H}$, that is,
$f^{\ast}=\operatorname*{\arg\min}_{g\in\mathbb{H}}\left\{\widehat{R}_{n}(g)+\lambda_{\textup{reg}}\|g\|_{\mathbb{H}}^{2}\right\},$ (6)
for some $\lambda_{\textup{reg}}>0$ lies on the span of the kernel sections generated by the sample. More explicitly, the solution of (6) has the form $f^{*}(\cdot)=\sum_{i=1}^{n}\alpha_{i}^{\ast}K(\cdot,\mathbf{z}_{i})$, where the vector $\bm{\alpha}^{\ast}=(\alpha_{1}^{\ast},\ldots,\alpha_{n}^{\ast})^{\top}\in\mathbb{R}^{n}$ can be determined by solving the regularized linear (Gramian) regression problem associated to the Gram matrix $\mathsf{K}$ of the sample, that is,
@@
where $\mathsf{Y}_{n}:=(y_{1},\ldots,y_{n})^{\top}$ and $\|\cdot\|_{2}$ denotes the Euclidean norm. The optimization problem (7) admits the closed form solution
$\bm{\alpha}^{*}=\left(\mathsf{K}^{\top}\mathsf{K}+\lambda_{\textup{reg}}\mathsf{K}\right)^{-1}\mathsf{K}\mathsf{Y}_{n}.$ (8)

### v3

@@ I Introduction
Despite the ease of training associated with the RC methodology, its implementation depends heavily on hyperparameters that are sometimes difficult to estimate robustly. This difficulty has motivated various authors to replace the standard RC approach with nonlinear vector autoregression in which the covariates are monomials constructed using the previous inputs. In these methods, the only hyperparameters to be chosen are the maximum order of the monomials and the number of lags of past signals. This approach has been called next-generation reservoir computing (NG-RC) [22, 23, 24, 25, 26, 27] and has displayed excellent performance in control and spatio-temporal forecasting tasks.
@@
The last part of the paper contains various numerical simulations that illustrate that (i) using the more computationally tractable polynomial kernel regression, one has access to a broader feature space, which allows learning more complex systems than those typically presented when using the NG-RC methodology and (ii) the Volterra kernel is a useful tool that can produce more accurate forecasts because it allows one to take infinite lags and monomial orders into account, which is of relevance in the modeling of long-memory phenomena that exhibit functionally complex dependencies. All the codes and data necessary to reproduce the results in the paper are available at [32].
@@ II.1 Notation
Let $\mathbb{Z}$ denote the set of integers and ${\mathbb{Z}_{-}}$ be the set of non-positive integers. Denote by $\mathbb{R}$ the set of real numbers and the $d$-dimensional Euclidean space by $\mathbb{R}^{d}$. We often work with sequence spaces. Let the space of sequences of $d$-dimensional real vectors indexed by ${\mathbb{Z}_{-}}$ be denoted $\left(\mathbb{R}^{d}\right)^{\mathbb{Z}_{-}}$. Given $\tau\in\mathbb{N}$, we shall use $\left(\mathbb{R}^{d}\right)^{\tau}$ for the space of $\tau$-long sequences of $d$-dimensional real inputs indexed by $\{-\tau+1,\ldots,-1,0\}$. In the description of sequence spaces, we may also replace $\mathbb{R}^{d}$ with any subset of finite-dimensional Euclidean space $\mathcal{W}$. The components of sequences $\underline{\mathbf{z}}\in\mathcal{W}^{\mathbb{Z}_{-}}$ are given by ${\bf z}_{i}\in\mathcal{W}$, $i\in\mathbb{Z}_{-}$, that is, $\underline{\mathbf{z}}=\left({\bf z}_{i}\right)_{i\in\mathcal{W}}$. Now, given $i\in\mathbb{Z}_{-}$, $\tau\in\mathbb{N}$, and $\mathbf{z}_{i}\in\mathcal{W}$, we denote by $\underline{\mathbf{z}_{i}}\coloneqq(\ldots,\mathbf{z}_{i-1},\mathbf{z}_{i})\in{\cal W}^{\mathbb{Z}_{-}}$ and by $\underline{\mathbf{z}_{i}}^{\tau}\coloneqq(\mathbf{z}_{i-\tau+1},\ldots,\mathbf{z}_{i})\in{\cal W}^{\tau}$. We use $\mathcal{Z}$ and $\mathcal{Y}$ to refer to the input space and to the output space, respectively. The examples of $\mathcal{Z}$ can be $\mathbb{R}^{d}$, $\left(\mathbb{R}^{d}\right)^{\tau}$, $\left(\mathbb{R}^{d}\right)^{\mathbb{Z}_{-}}$, $\mathcal{W}^{\mathbb{Z}_{-}}$, etc. The output space $\mathcal{Y}$ is typically finite-dimensional subsets of Euclidean space, or even subsets of $\mathbb{R}$.
@@ II.2 Data generating processes
Throughout this paper, we are interested in the learning and forecasting of discrete-time, time-invariant, and causal dynamic processes that are generated by functionals of the form $H:\left(\mathbb{R}^{d}\right)^{\mathbb{Z}_{-}}\longrightarrow\mathcal{Y}$, that is, given $\underline{{\bf z}}\in({\mathbb{R}}^{d})^{\mathbb{Z}}$, we have that ${\bf y}_{t}=H(\underline{{\bf z}_{t}})$, $t\in\mathbb{Z}$. In practice, the inputs $\underline{{\bf z}}\in({\mathbb{R}}^{d})^{\mathbb{Z}}$ can be just deterministic sequences or realizations of a stochastic process. A related data generating process that we shall be using are the causal chains with infinite memory (CCIM) [33, 34, 35]. These are infinite sequences $(\mathbf{y}_{i})_{i\in\mathbb{Z}}$, where $\mathbf{y}_{i}\in\mathcal{Y}$ for all $i\in\mathbb{Z}$, are such that $\mathbf{y}_{i}=H(\underline{\mathbf{y}_{i-1}})$ for all $i\in\mathbb{Z}$ and some functional $H:{\cal Y}^{\mathbb{Z}_{-}}\rightarrow{\cal Y}$. Takens’ Theorem [36] and its generalizations [37, 38, 39, 40] guarantee that the low-dimensional observations of dynamical systems follow, under certain conditions, a dynamical prescription of this type.
@@ II.3 Kernel methods
Kernels and induced RKHS. A kernel on $\mathcal{Z}$ is a function $K:\mathcal{Z}\times\mathcal{Z}\longrightarrow\mathbb{R}$ that is symmetric and positive semi-definite. In this context, positive semi-definiteness means that for any $\alpha_{i},\alpha_{j}\in\mathbb{R}$, $\mathbf{z}_{i},\mathbf{z}_{j}\in\mathcal{Z}$, $i,j\in\{1,\ldots,n\}$,
@@
Let $\mathbb{H}$ be a Hilbert space of real-valued functions on $\mathcal{Z}$ endowed with pointwise sum and scalar multiplication. Denote the inner product on $\mathbb{H}$ by $\langle\cdot,\cdot\rangle_{\mathbb{H}}$. We say that $\mathbb{H}$ is a reproducing kernel Hilbert space (RKHS) associated to the kernel $K$ if the following two conditions hold. First, for all $\mathbf{z}\in\mathcal{Z}$, we have that the functions $K(\cdot,\mathbf{z})\in\mathbb{H}$, and for all $\mathbf{z}\in\mathcal{Z}$ and all $f\in\mathbb{H}$, the reproducing property, $f(\mathbf{z})=\langle f,K(\cdot,\mathbf{z})\rangle_{\mathbb{H}}$, $z\in{\cal Z}$, is satisfied. The maps of the form $K_{\mathbf{z}}(\cdot)\coloneqq K(\cdot,\mathbf{z}):\mathcal{Z}\longrightarrow\mathbb{R}$ are called kernel sections. Second, the Dirac functionals $\delta_{{\bf z}}:\mathbb{H}\rightarrow\mathbb{R}$ defined by $\delta_{{\bf z}}(f):=f({\bf z})$ are continuous, for all ${\bf z}\in{\cal Z},$ with respect to the metric on $\mathbb{H}$ induced by the inner product $\langle\cdot,\cdot\rangle_{\mathbb{H}}$.
Canonical feature maps and spaces. In this setup one may define the map $\Phi_{\mathbb{H}}:\mathcal{Z}\longrightarrow\mathbb{H}$ given by $\Phi_{\mathbb{H}}(\mathbf{z})\coloneqq K_{\mathbf{z}}$, ${\bf z}\in{\cal Z}$. Then from the reproducing property of the RKHS, the kernel function can be written as the inner product of kernel sections. Indeed, given ${\bf z},{\bf z}^{\prime}\in{\cal Z}$. we can write
@@
where the bar denotes the completion with respect to the metric induced by the inner product $\langle\cdot,\cdot\rangle_{\mathbb{H}}$. This inner product obviously satisfies that for $f=\sum^{m}_{i=1}\alpha_{i}K_{\mathbf{z}^{\prime}_{i}},g=\sum^{n}_{j=1}\beta_{j}K_{\mathbf{z}_{j}}\in\mathbb{H}$, $\langle f,g\rangle_{\mathbb{H}}=\sum_{i=1}^{m}\sum_{i,j=1}^{n}\alpha_{i}\beta_{j}K(\mathbf{z}^{\prime}_{i},\mathbf{z}_{j})$. We use the symbol $\|\cdot\|_{\mathbb{H}}$ for the norm induced by $\langle\cdot,\cdot\rangle_{\mathbb{H}}$.
Feature maps and their associated kernels. The construction we just presented produces an RKHS and a feature map out of a kernel map. Conversely, given any Hilbert space $(H,\left\langle\cdot,\cdot\right\rangle_{H})$ and a map $\Phi:\mathcal{Z}\rightarrow H$ from the set ${\cal Z}$ into $H$, we can construct a kernel map that has $\Phi$ as a feature map and $H$ as a feature space, that is, define
@@
Such a function $K$ is clearly symmetric and positive semi-definite; thus, it satisfies the conditions needed to be a kernel function. It can actually be proved (see [30, Theorem 4.16]) that a map $K:\mathcal{Z}\times\mathcal{Z}\longrightarrow\mathbb{R}$ is a kernel if and only if there exists a feature map that allows $K$ to be represented as in (1). Note that given a kernel function, the feature representation (1) is not unique in the sense that neither the feature space $H$ nor the feature map $\Phi:\mathcal{Z}\rightarrow H$ are unique. In particular, given $K$, the canonical feature map and the RKHS are a choice of feature map and feature space for $K$, respectively.
@@
=$\mathbb{H}=\{f:\mathcal{Z}\rightarrow\mathbb{R}\mid f(\cdot)=\left<w,\Phi(\cdot)\right>_{H},w\in H\},$ (2)
@@
Kernel ridge regression. Suppose that ${\cal Z}\subset{\mathbb{R}}^{d}$, $\mathcal{Y}\subset\mathbb{R}$, and that a function $f:{\cal Z}\rightarrow{\cal Y}$ needs to be estimated using a finite sample of $n$ pairs of input and outputs $\{(\mathbf{z}_{i},y_{i}:=f(\mathbf{z}_{i}))\}_{i=1}^{n}$ where $\mathbf{z}_{i}\in\mathcal{Z}$ and $y_{i}\in\mathcal{Y}$, $i=1,\ldots,n$. If the estimation is carried out using an empirical risk with squared loss
@@
and the hypothesis set is the RKHS $\mathbb{H}$ corresponding to a kernel $K$, the Representer Theorem [29] states that the minimizer of the ridge regularized empirical risk on $\mathbb{H}$, that is,
$f^{\ast}=\operatorname*{\arg\min}_{g\in\mathbb{H}}\left\{\widehat{R}_{n}(g)+\Omega(\|g\|_{\mathbb{H}}^{2})\right\},$ (4)
where $\Omega:(0,\infty)\longrightarrow\mathbb{R}$ is any increasing function, lies on the span of the kernel sections generated by the sample. More explicitly, the solution of (4) has the form
$f^{*}(\cdot)=\sum_{i=1}^{n}\alpha_{i}^{\ast}K(\cdot,\mathbf{z}_{i}),$ (5)
where the vector $\bm{\alpha}^{\ast}=(\alpha_{1}^{\ast},\ldots,\alpha_{n}^{\ast})^{\top}\in\mathbb{R}^{n}$ can be determined by solving the regularized linear (Gramian) regression problem (nonlinear in inputs, linear in covariates) associated to the Gram matrix $\mathsf{K}$ of the sample, that is,
$\min_{\bm{\alpha}\in\mathbb{R}^{n}}\left\{\frac{1}{n}\|\mathsf{K}\bm{\alpha}-\mathsf{Y}_{n}\|_{2}^{2}+\Omega(\bm{\alpha}^{\top}\mathsf{K}\bm{\alpha})\right\},$ (6)
where $\mathsf{Y}_{n}:=(y_{1},\ldots,y_{n})^{\top}$ and $\|\cdot\|_{2}$ denotes the Euclidean norm. Let the regularization function $\Omega:(0,\infty)\longrightarrow\mathbb{R}$ be defined as $\Omega(u)=\lambda_{\textup{reg}}u$ for some regularization strength $\lambda_{\textup{reg}}>0$. Then the optimization problem (6) can be rewritten as
@@
which admits the closed form solution
$\bm{\alpha}^{*}=\left(\mathsf{K}^{2}+\lambda_{\textup{reg}}\mathsf{K}\right)^{-1}\mathsf{K}\mathsf{Y}_{n}.$ (8)
@@
[…]
This time around, the bar denotes the uniform closure. Note that the continuity of $K$, the reproducing property, and the compactness of ${\mathcal{K}}$ imply that the uniform closure that defines $K\left({\mathcal{K}}\right)$ contains the completion of the vector space ${\rm span}\left\{K_{\mathbf{z}}\mid\mathbf{z}\in{\cal K}\right\}$. The continuous kernel $K$ is called kernel universal when for any compact subset $\mathcal{K}\subset\mathcal{Z}$,
$K(\mathcal{K})=C^{0}(\mathcal{K}).$
Many kernels used in practice are universal. When a kernel is universal, the RKHS associated to the kernel, approximates arbitrarily well all continuous functions defined on compacta $\mathcal{K}$. This implies that having kernel universality ensures that the corresponding RKHS used in the kernel ridge regression problems discussed in the previous section is rich enough to approximate arbitrarily well any target (continuous) function. Examples of universal kernels include the Gaussian kernel and the Volterra kernel [31], which we discuss in detail later on in Section IV. On the other hand, the polynomial kernel, although popular, is not universal [44].